OpenStudy (anonymous):

What is the slope of this line? y= 2/3x +3 a) 3 b) 2 c) 2/3 Second question: Could this ordered pair be a sloution to this equation? y=2x-1 (0,-1) a) no b) yes

OpenStudy (btaylor):

1. In a line of form \(y=m \times x + b\) the slope is \(m\). 2. Plug in 0 for x and -1 for y. If it comes out to something like 2=2, then it is a solution.

OpenStudy (yttrium):

always remember this. \[y = mx+b\] such that m is the slope. inthe second question, just substitute the values for the liner equation (0,-1) -1 = (2/3)(0)-1 -1 = -1 hence, true :) (means the orderred pair is applicable)

OpenStudy (anonymous):

Thank you for the quick answers, it was urgent!

OpenStudy (yttrium):

my three medals were lost :((

OpenStudy (anonymous):

Why? :o